OpenStudy (anonymous):

Simplify 6/5-sqrt3 The possible answers are a.6sqrt3/5sqrt3-3 b.30+sqrt3/22 c.6sqrt3/2 d.15+3sqrt3/11

OpenStudy (jdoe0001):

\(\bf \cfrac{6}{5-\sqrt{3}}?\)

OpenStudy (math&ing001):

Multiply the numerator and the denominator by (5+sqrt(3)) to remove the root from the denominator : \[\frac{ 6 }{ 5-\sqrt{3} }=\frac{ 6 }{ 5-\sqrt{3} }*\frac{ 5+\sqrt{3} }{ 5+\sqrt{3} }\] Simplify it to get your answer.

OpenStudy (anonymous):

is it 15 +3sqrt3/11

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

Thanks! @math&ing001

OpenStudy (math&ing001):

You're welcome !

OpenStudy (math&ing001):

And welcome to Openstudy ! If you need any help about how to use the website message me.

OpenStudy (anonymous):

okay great! @math&ing001